OpenStudy (anonymous):

Use cross products to demonstrate whether or not the two ratios are equivalent: 4/6 and 14/20.

OpenStudy (alekos):

just convert the fractions to the lowest common denominator

OpenStudy (texaschic101):

cross products... 4/6 = 14/20 (4)(20) = (6)(14) 80 = 84.....not equivalent

OpenStudy (anonymous):

ohh okay thank you so much

OpenStudy (texaschic101):

no problem :)